Write the rate law for the two elementary reaction equations. 2 A(g) + B(g) → C(g) +D(8) ratek X(g) + Y() - (8) rate =k
Answer #1

Rate laws or rate equations are mathematical expressions that describe the relationship between the rate of a chemical reaction and the concentration of its reactants. In general, a rate law (or differential rate law, as it is sometimes called) takes this form:

rate=k[A]m[B]n[C]p

in which [A], [B], and [C] represent the molar concentrations of reactants, and k is the rate constant, which is specific for a particular reaction at a particular temperature.The exponents m, n, and p are usually positive integers (although it is possible for them to be fractions or negative numbers).

for first equation,

2A + B = C + D

rate= k [A]2 [B]

for second equation,

X + Y = Z

rate= k [X] [Y]